OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 14, 22 …, if there are 26 terms? 2,754 2,756 2,758 2,760

OpenStudy (callisto):

1. Find the last term using \[a_n = a + (n-1)d\] where a is the first term , n is the number of term, d is the common difference. Common difference = 14-6 = 8 Last term = 6 + (26-1) (8) =? 2. Find the sum using the formula \[S_n = \frac{n}{2}(a+a_n)\], where n is the number of term. \[Sum =\frac{ 26}{2}(6+?) =...\] Can you do it?

OpenStudy (anonymous):

no i dont know this but im on a timed exam that decides my graduation

OpenStudy (callisto):

All you need to do is to use your calculator.. First, find the last term using the calculator as I've given you the formula already (with all the numbers you need:S)

OpenStudy (anonymous):

i dont have a calculator

OpenStudy (callisto):

... Use the calculator on your pc

OpenStudy (anonymous):

i dont know that i have that or even where to find it

OpenStudy (callisto):

Use this then: http://www.wolframalpha.com/

OpenStudy (callisto):

Type (or copy) the formula, it will give you the answer

OpenStudy (amistre64):

im not sure if this is the right place to be doing your exams if you dont know the material yet.